OpenStudy (khameleon):

If you multiplied a number by 0.5, the result would be A) twice the original number. B) half of the original number. C) a fifth of the original number. D) five times the original number.

KjSaif (kjsaif):

Here's a few examples

KjSaif (kjsaif):

2* 0.5 = 1

OpenStudy (will.h):

Lool I was gonna write that same exact example

KjSaif (kjsaif):

5 * 0.5 = 2.5

KjSaif (kjsaif):

lol @Will.H

KjSaif (kjsaif):

10 * 0.5 = 5

OpenStudy (will.h):

Good work mate

KjSaif (kjsaif):

so what do you think would be the right answer @khameleon

OpenStudy (jesselacompte):

10 * 0.5 = 5

KjSaif (kjsaif):

were the answers Twice the amount? Half the amount? a fifth of the amount? or fife times the amount?

KjSaif (kjsaif):

amount = original number

OpenStudy (will.h):

He's thinking.. Give em a break

OpenStudy (khameleon):

c

OpenStudy (will.h):

Well...

OpenStudy (jesselacompte):

c

KjSaif (kjsaif):

hmmmmm not quite................

OpenStudy (will.h):

Good luck Saif

KjSaif (kjsaif):

think about it......was it half the amount? or fifth the amount

KjSaif (kjsaif):

half is this.................2-----> 1

OpenStudy (will.h):

Dude 0.5 means half

OpenStudy (will.h):

Fifth means 1/5

KjSaif (kjsaif):

^^

KjSaif (kjsaif):

so looking at it again........What do you think?

KjSaif (kjsaif):

2 x 0.5 = 1

KjSaif (kjsaif):

hmmmm...........thats half of 2

KjSaif (kjsaif):

So the answer is....................

KjSaif (kjsaif):

what do you think @khameleon

KjSaif (kjsaif):

Do you think it's half?

KjSaif (kjsaif):

or no?

KjSaif (kjsaif):

Yep The answer is half of the original number.